A recent study shows that the annual cost of maintaining a building in Accra averages 200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items associated with the maintenance of building i.e. everything is made 30% more expensive. Calculate the standard deviation of the annual cost of maintaining a building in Accra
1
Expert's answer
2020-05-12T18:32:58-0400

"\text {Let X represents the annual cost of maintaining}\\
\text { a building in Accra, then, E(x)=200, V(x)=260,}\\
\text {Let Y represents the annual cost of maintaining}\\
\text { a building in Accra after adding a tax of 30\%,}\\
Y=1.3X\\
V(Y)=V(1.3X)=1.3^{2}V(x)\\
=1.3^{2}(260)=439.4,\\
\therefore \sigma_{Y}=\sqrt{V(Y)}=\sqrt{439.4}=20.962"